matthew's bakery sells strawberry cakes for $15 each. Matthews variable costs are $5 and her fixed costs total $3780 per month. if matthew's tax rate is 25% how many cakes mus matthew sell each month if she wants to earn $7560 in net income?
Answer #1

Income before tax = 7560*100/75 = 10080

Required cakes = (Fixed cost+Net income)/Contribution margin per unit

                          = (3780+10080)/10

Required cakes = 1386 Cakes